# EcoServe: Designing Carbon-Aware AI Inference Systems

### v1

@@ Abstract
The rapid increase in LLM ubiquity and scale levies unprecedented demands on computing infrastructure. These demands not only incur large compute and memory resources, but also significant energy, yielding large operational and embodied carbon emissions. In this work, we present two main observations. First, while GPUs dominate operational carbon, host processing systems (e.g., CPUs, memory, storage) dominate embodied carbon. Second, based on traces from production deployment of two Generative AI services in the cloud, offline, batch-inference accounts for a significant portion (up to 50%) of serving capacity. We propose four pillars of carbon-conscious infrastructure design for LLM serving systems: Reduce, Reuse, Rightsize, and Recycle. We demonstrate that EcoServe can lower carbon emissions by up to 47%, compared to performance, energy, and cost-optimized design points, while maintaining performance targets and SLOs.
I Introduction
@@
The rapid growth of Large Language Models (LLMs) has led to unprecedented demands on computing infrastructure. From 2019 to 2022, we have seen an over 100$\times$ increase in model size from GPT-2 (1.5 billion parameters) to GPT-3 (175 billion parameters) to PaLM (540 billion parameters) [12]. The increasing compute demands have also led to dramatic increases in the power and energy demands for deploying LLM’s at-scale [67, 44]. Given the increasing demand and corresponding energy consumption, the Electric Power Research Institute (EPRI) projects that AI-driven data centers will account for up to 40% of the power consumed in local grids (i.e. Virgina, Oregon) [2]. The rising compute demand of deploying LLMs not only impacts the efficiency of the data center, but also the overall environmental impact of GenerativeAI and data centers. The increasing power and energy demands for LLM inference have led to significant research from the machine learning, systems, and computer architecture community. Systems such as vLLM [68], SGLang [75], Sarathi-serve [6], DynamoLLM [62], Melange [38], and Splitwise [53] have made significant strides in optimizing GPU utilization, throughput, cost, and energy efficiency via various intelligent resource provisioning, allocation, and scheduling strategies.
Figure 1: (Left) Breakdown of power (TDP) and embodied carbon between host systems (CPU) and GPU. (Right) EcoServe optimizes operational and embodied carbon based on input hardware, model, and workload traces by reusing CPU for offline inference, rightsizing GPUs, reducing underutilized hardware, and recycling hardware.
Although current approaches to LLM serving optimization focus on performance, power, energy efficiency, and cost, they do not directly address the growing environmental impact of LLM deployment. Crucially, recent work shows that the carbon footprint of AI is determined not only by operational emissions but also by embodied emissions [31, 71, 70]. While operational emissions are a product of energy consumption and carbon intensity of local power grids, embodied emissions owe to carbon impact from hardware manufacturing (i.e., application processors, memory, storage) data center construction. Recent academic and industry efforts to understand the environmental impact of data centers, as well as publicly available industrial sustainability reports show that more than 50% carbon emissions from hyper-scalars driven by renewable energy are due to embodied carbon [31, 39, 47, 27, 45]. Although significant attention has been paid to operational and energy efficiency, the embodied carbon footprint of hardware components remains largely overlooked.
Embodied carbon poses new trade-offs in the design space. Figure 1 illustrates the power and embodied carbon breakdown of two AI inference systems comprising NVIDIA A100 and H100 GPUs. While the GPUs account for around 75% of the thermal dissipated power (TDP), the host-processing system accounts for around 75% of embodied carbon. The high embodied overheads from host systems owe to a combination of memory capacity, storage, and mother-board (e.g., PCB, peripheral interconnects). While GPUs dominate the power consumption of AI servers, the embodied carbon is dominated by the host systems. Furthermore, during LLM inference, despite the CPUs being responsible for key operations like tokenization, graph computation, tensor management, and scheduling, they exhibit poor utilization; we observe around 6% average utilization per core based on characterization with vLLM (v0.6.2) on a 16-core server, effectively wasting available resources. Finally, embodied carbon of newer AI inference systems grows as they implement more compute, memory bandwidth, memory capacity, and storage.
In order to tackle these challenges, in this paper we present EcoServe, a carbon-aware resource allocation and scheduling framework for LLM serving. Central to EcoServe is co-designing decisions across capacity planning to minimize embodied carbon, resource allocation based on workload and heterogeneous hardware, and runtime scheduling to minimize operational emissions. We find separately considering these phases leads to sub-optimal outcomes i.e., improving energy efficiency can come at the expense of reducing overall carbon emissions. This misalignment is particularly pronounced for LLM serving workloads, where the compute-bound prefill phase and memory-bound decode phase demand different resources. To guide carbon-aware AI serving system design, we first develop an embodied carbon model, built upon existing models for individual integrated circuits [30, 24, 34], tailored for AI systems by considering the full host system (e.g., memory, storage, PCB, and peripheral circuits) and GPUs (e.g., memory, GPU die, peripheral ICs). We identify Host CPU as an untapped opportunity, and argue that optimizing the carbon footprint requires leveraging the capacity of CPU inference.
Using the model, EcoServe’s cross-layer approach focuses on four key sustainability-driven design principles: a) reusing existing CPU and host memory systems to increase AI cluster capacity for offline inference which accounts for nearly 30% of inference capacity based on traces from a production cloud provider, b) right-sizing host and accelerator systems for various model size and phases, and input sizes to balance performance, efficiency, and carbon, with hardware-software co-design, c) reducing wasted embodied carbon by eliminating unused resources for leaner hardware at no performance overheads, and finally, d) recycling host systems across generations of AI servers while still replacing accelerators at the current rate, to optimize collectively for efficiency and carbon. We quantify and demonstrate these solutions across a wide array of models and deployment scenarios with 1) open-source traces like ShareGPT [13] and Azure Function Traces [53]; 2) traces from a cloud service provider for two production LLM applications.
@@
  • We introduce EcoServe, a carbon-aware framework for LLM infrastructure design that optimizes both operational and embodied emissions. Fundamentally EcoServe comprises four design strategies—Reuse, Rightsize, Reduce, and Recycle—motivated by insights from fine-grained embodied carbon models for AI systems and detailed profiling and analysis of inference performance and efficiency across diverse hardware and serving environments (e.g., LLMs, SLO targets, demand load).
@@
  • Putting the EcoServe optimizations together using an ILP formulation to co-optimize performance, efficiency, and carbon, we demonstrate 1.4x-2.2 $\times$ carbon benefits with minimal performance degradation for both online and offline workloads across a variety of LLM models and workloads.
II Background
@@
SoC’s Tech. & Area dependent ACT [30], iMec [24]
DDR4/LPDDR5 0.29 / GB TechInsights [35]
GDDR6 0.36 / GB TechInsights [35]
HBM2 0.28 / GB TechInsights [35]
HBM3e 0.24 / GB TechInsights [35]
SSD 0.110 / GB R740 LCA [10] SCARIF  [34]
PCB 0.048 / cm2 12 layers Dell R740 LCA [10]
Ethernet card 4.91 Dell R740 LCA [10]
HDD controller 5.136 Dell R740 LCA [10]
Cooling 7.877 per 100W Scaled with TDP[34]
PDN / PSU 3.27 per 100W Scaled with TDP, Schneider [59]
TABLE I: Embodied carbon modeling strategy for each hardware component in AI inference systems.
Carbon Emissions in Computing Infrastructure Computing infrastructure contributes to carbon emissions in two primary ways: operational carbon and embodied carbon. Operational carbon refers to emissions from energy consumption during system operation, including direct power consumption by computing components and indirect emissions from cooling systems.
Embodied carbon, often overlooked, encompasses emissions from manufacturing, transportation, and eventual disposal of hardware components. Prior works indicate that for typical data center hardware, embodied carbon can represent over 50% of total lifetime emissions, with memory and storage being particularly carbon-intensive in their manufacturing process [69, 64, 31].
Recent work like ACT, SCARIF, FOCAL, LLMCarbon [30, 34, 22, 21] has highlighted the importance of considering both types of emissions. However, existing modeling approaches often do not look at fine-grained GPU carbon vs Host carbon footprint and treat these as separate concerns rather than considering their complex interactions with workloads in real-world systems.
LLM Serving Systems
State-of-the-art LLM serving systems like vLLM, TensorRT-LLM, AlpaServe, SGLang, Sarathi-serve, DistServe, and DeepSpeed [55, 38, 6, 76, 53, 75, 4, 41] have made significant advances in optimizing resource utilization LLM inference.
@@
To address these challenges, these inference systems leverage GPU hybrid parallelization strategies, combining data, tensor, and pipeline parallelism to optimize computation across GPUs. Advanced memory management techniques, such as activation checkpointing, memory offloading, and paged attention, help alleviate memory pressure and improve resource utilization. Scheduler optimizations, including dynamic batching and chunked prefill processing, and disaggregating prompting and decoding phases, ensure high throughput while meeting latency constraints, enabling these systems to efficiently handle diverse LLM inference workloads [53].
@@
While these systems primarily focus on performance metrics such as throughput and latency, recent works like DynamoLLM [62] and Melange [28] have brought attention to energy and cost efficiency of LLM inference. However, optimizing for energy consumption or cost does not directly optimize carbon impact.
III Understanding AI Systems’ Carbon Footprint
@@
Figure 2: Trends in bit density (left) and embodied carbon footprint (right) across varoius DRAM memory technologies for 3 different manufacturers.
@@
Based on recent efforts quantifying the carbon footprint of computing systems [30, 34, 21], the aggregate carbon emissions of AI systems is the combination of operational and embodied carbon as shown in the equation below:
$CF_{Task}=(P_{host}+P_{gpu})\times t\times CI+CF^{emb}_{host}\frac{t}{LT}+CF^{emb}_{gpu}\frac{t}{LT}$
@@
where $P$ represents the power consumption of the host and GPU, $t$ represents the workload run-time, $LT$ represents the hardware lifetime, and $CF^{emb}$ represents the host and GPU embodied carbon. In terms of operational emissions, we measure the power consumption of host-processing systems including CPUs, DRAM, SSD, using a combination of Intel RAPL [17]; the power consumption of GPUs is empirically measured using NVML [49]. The carbon intensity of the different data center power grids is based on geo-temporal variance of power grid carbon intensities [5, 69]. To estimate embodied carbon, we leverage a combination of openly available product environmental reports and life cycle analyses from industry [10] and recently published, open-source tools to quantify the emissions from hardware manufacturing including the Architectural Carbon Modeling Tool (ACT), SCARIF, and GreenSKU [30, 34, 69]. In this section we describe how we extend current methods to study the embodied carbon of AI inference platforms including host systems, motherboards, and GPUs.
III-A Modeling Embodied Carbon of AI Systems
The embodied carbon footprint (CF) owes to a combination of components for the host systems and GPUs, as shown below:
$CF^{emb}_{host,gpu}=\sum_{k}^{\text{DRAM, PWB, PDN, Cooling, Storage, SoC Die,Chasis}}CF_{k}$
@@ III-A Modeling Embodied Carbon of AI Systems
For host systems, we must account for the main CPU die, DRAM memory, SSD and HDD storage, the main printed wiring board (PWB), power distribution network (PDN), and server chassis. Similarly, for the GPUs we must account for the main SoC, DRAM memory, PDN, cooling (e.g., heat sink), and PWB. Table I describes the salient parameters in the embodied carbon model for AI inference systems. For many components including PDN and cooling (e.g., heat sink) given the lack of specific technology information we rely on publicly available product environmental reports from Dell [10]. Below we detail the methodology for estimating the embodied carbon of the remaining salient components.
III-A1 Integrated circuits (Application Processors)
@@
To compute the embodied carbon emissions of application processors (e.g., CPU die, GPU die) we rely on recently published architectural carbon modeling tools [30, 34, 24, 21]. For application processors we input the process technology and die area into ACT [30, 24].
III-A2 Memory and Storage

### v2

@@ Abstract.
The rapid increase in LLM ubiquity and scale levies unprecedented demands on computing infrastructure. These demands not only incur large compute and memory resources but also significant energy, yielding large operational and embodied carbon emissions. In this work, we present three main observations based on modeling and traces from the production deployment of two Generative AI services in a major cloud service provider. First, while GPUs dominate operational carbon, host processing systems (e.g., CPUs, memory, storage) dominate embodied carbon. Second, offline, batch inference accounts for a significant portion (up to 55%) of serving capacity. Third, there are different levels of heterogeneity across hardware and workloads for LLM inference. Based on these observations, we design EcoServe, a carbon-aware resource provision and scheduling framework for LLM serving systems. It is based on four principles - Reduce, Reuse, Rightsize, and Recycle (4R). With a cross-stack ILP formulation and design, we demonstrate that EcoServe can lower carbon emissions by up to 47%, compared to performance, energy, and cost-optimized design points, while maintaining performance targets and SLOs.
copyright: none
1. Introduction
@@
The rapid growth of Large Language Models (LLMs) has led to unprecedented demands on computing infrastructure. From 2019 to 2022, we have seen an over 100$\times$ increase in model size from GPT-2 (1.5 billion parameters) to GPT-3 (175 billion parameters) to PaLM (540 billion parameters) (chowdhery2023palm, ). The increasing compute demands have also led to dramatic increases in the power and energy demands for deploying LLM’s at-scale (tschand2024mlperf, ; luccioni2024power, ). The Electric Power Research Institute (EPRI) projects that AI-driven data centers will account for up to 40% of the power in local grids (i.e. Virgina, Oregon) (epri2024, ).
The rising compute demand of deploying LLMs impacts not only the efficiency of the data center but also the overall environmental impact of GenerativeAI and data centers.
Figure 1. (Left) Breakdown of thermal design power provision (TDP) and embodied carbon between host systems (CPU) and GPU. (Right) Through the 4R strategy, EcoServe optimizes the carbon savings for various AI workloads.
The systems community has responded with numerous innovations to optimize LLM inference. Systems like vLLM (vllm2024perfupdate, ), SGLang (zheng2023efficiently, ), and DynamoLLM (stojkovic2024dynamollm, ) have significantly improved GPU utilization, throughput, and energy efficiency via sophisticated resource management. However, these approaches optimize primarily for operational metrics, overlooking a crucial aspect of AI’s environmental impact: embodied carbon emissions from hardware manufacturing.
Recent sustainability reports from hyperscalers reveal that embodied carbon now accounts for over 50% of their total emissions when using renewable energy (gupta2021chasing, ; microsoftSustainabilityReport, ; GoogleSustainabilityReport, ). This presents a fundamental challenge: optimizing operational efficiency alone may not minimize total carbon impact. Our analysis reveals a critical insight: while GPUs dominate operational power consumption, CPUs and host systems contribute disproportionately to embodied carbon (Figure 1). Note that TDP can serve as a proxy for operational carbon assuming a similar high utilization workload profile and same carbon intensity (assuming 80% for both CPU and GPU). The high embodied overheads from host systems owe to a combination of memory capacity, storage, and mother-board (e.g., PCB, peripheral interconnects). This misalignment creates an opportunity to rethink how we design AI infrastructure.
Furthermore, AI inference workloads exhibit workload and hardware heterogeneity—differences in model size, prompt length, decoding patterns, workload phases and SLO requirements—that create new optimization opportunities. However, existing approaches fail to leverage all the heterogeneity dimensions, and co-optimize capacity planning, resource allocation, and runtime scheduling, leading to misalignment between provisioned resources and runtime carbon reduction. For example, a strategy that only optimizes operational efficiency may increase embodied carbon by requiring frequent hardware upgrades or over-provisioning GPU capacity.
To holistically optimize embodied and operational emissions, we introduce EcoServe, a carbon-aware AI inference framework that co-designs capacity planning, resource allocation, and scheduling. Unlike prior work that treats these phases independently, EcoServe’s cross-layer design ensures that capacity planning decisions are effectively exploited at runtime. Specifically, while capacity planning determines long-term provisioning to reduce embodied carbon, runtime scheduling dynamically adapts these provisions to workload variation, ensuring optimal hardware utilization without increasing embodied or operational overhead.
At the core of EcoServe is a sustainability-driven optimization framework based on four key principles: Reuse, Rightsize, Reduce, and Recycle (4R). These principles collectively minimize carbon impact across different dimensions of system operation:
  • Reuse: Exploits underutilized host CPUs for offline inference to amortize embodied carbon across workload phases.
  • Rightsize: Dynamically provisions GPUs and CPUs based on model size, execution phase, and workload demand to avoid over-provisioning.
  • Reduce: Eliminates unnecessary host memory and storage resources to reduce embodied carbon overhead.
  • Recycle: Extends the lifetime of host processing systems while selectively upgrading accelerators to balance embodied and operational trade-offs.
Figure 1(right) shows the impact of each of these design strategies on two scenarios: offline or batch processing, and online serving. The individual strategies achieve carbon savings of 29-41% compared to performance-optimized baselines. While reduce and right size benefit carbon savings in online serving, offline serving benefits from a combination of reuse, reduce, and rightsize. Altogether the different design strategies save up to 2$\times$ total carbon for both offline and online serving. More generally, we evaluate EcoServe using both open-source datasets (ShareGPT (sharegpt, ), Azure Functions (patel2023splitwise, )) and production traces from a major cloud provider, as well as across heterogeneous model and hardware setups. Our results demonstrate EcoServe’s co-design solutions yield 1.4-2.2$\times$ reduction in total carbon emissions while maintaining performance targets across diverse LLM deployment scenarios.
@@
  • We introduce EcoServe, a carbon-aware framework for LLM infrastructure design that optimizes both operational and embodied emissions. Fundamentally EcoServe comprises four design strategies—Reuse, Rightsize, Reduce, and Recycle—motivated by insights from fine-grained embodied carbon models for AI systems and detailed profiling and analysis of inference performance and efficiency across diverse hardware and serving environments.
@@
  • Putting the EcoServe optimizations together using an ILP formulation to co-optimize performance, efficiency, and carbon, we demonstrate 1.4-2.2 $\times$ carbon benefits with minimal performance degradation for both online and offline workloads across a variety of LLM models and workloads.
By framing AI inference sustainability as a co-design problem across system stacks, EcoServe represents a paradigm shift towards carbon-aware systems design. Our results show that carbon-efficient AI infrastructure need not come at the cost of performance—by leveraging workload and hardware heterogeneity, sustainable AI inference can be both high-quality and performant.
2. Background
@@
SoC’s Tech. & Area dependent ACT (ACT, ), iMec (imec-dtco20, )
DDR4/LPDDR5 0.29 / GB TechInsights (TechInsights, )
GDDR6 0.36 / GB TechInsights (TechInsights, )
HBM2 0.28 / GB TechInsights (TechInsights, )
HBM3e 0.24 / GB TechInsights (TechInsights, )
SSD 0.110 / GB Dell R740 LCA (dellr740, ) + SCARIF (SCARIF, )
PCB 0.048 / cm2 12 layers Dell R740 LCA (dellr740, )
Ethernet card 4.91 Dell R740 LCA (dellr740, )
HDD controller 5.136 Dell R740 LCA (dellr740, )
Cooling 7.877 per 100W Scaled with TDP(SCARIF, )
PDN / PSU 3.27 per 100W Scaled with TDP, Schneider (se_lifecycle_calculator, )
Table 1. Embodied carbon modeling strategy for each hardware component in AI inference systems.
Figure 2. EcoServe’s carbon modeling framework with more fine-grained embodied carbon estimation on memory, storage, and power-related components. We highlighted the differences with ACT in the red box.
Carbon Emissions in Computing Infrastructure. Computing infrastructure contributes to carbon emissions in two primary ways: operational carbon and embodied carbon. Operational carbon refers to emissions from energy consumption during system operation, including direct power consumption by computing components and indirect emissions from cooling systems.
Embodied carbon, often overlooked, encompasses emissions from manufacturing, transportation, and eventual disposal of hardware components. Prior works indicate that for typical data center hardware, embodied carbon can represent over 50% of total lifetime emissions, with memory and storage being particularly carbon-intensive in their manufacturing process (greensku, ; dirty_ssd, ; gupta2021chasing, ).
Recent work like ACT, SCARIF, FOCAL, LLMCarbon (ACT, ; SCARIF, ; faiz2023llmcarbon, ; eeckhout2024focal, ) has highlighted the importance of considering both types of emissions. However, existing modeling approaches often do not look at fine-grained GPU carbon vs Host carbon footprints. We extend these works by providing a more fine-grained carbon modeling framework, especially for GPU systems, and show how these models can be used to optimize carbon footprint of LLMs .
LLM Serving Systems. State-of-the-art LLM serving systems like vLLM, TensorRT-LLM, AlpaServe, SGLang, Sarathi-serve, DistServe, and DeepSpeed (Deepspeed, ; kwon2023efficient, ; agrawal2024taming, ; zhong2024distserve, ; patel2023splitwise, ; zheng2023efficiently, ; TensorRT, ; alpaserve, ) have made significant advances in optimizing resource utilization LLM inference.
@@
To address these challenges, these systems leverage GPU hybrid parallelization strategies, combining data, tensor, and pipeline parallelism to optimize computation across GPUs (mei2024helix, ; li2022amp, ). Advanced memory management techniques, such as activation checkpointing, memory offloading, and paged attention, help alleviate memory pressure and improve resource utilization. Scheduler optimizations, including dynamic batching and chunked prefill processing, and disaggregating prompting and decoding phases, ensure high throughput while meeting latency constraints, enabling these systems to efficiently handle diverse LLM inference workloads (patel2023splitwise, ; zhong2024distserve, ).
@@
While these systems primarily focus on performance metrics such as throughput and latency, recent works like DynamoLLM (stojkovic2024dynamollm, ) and Melange (griggs2024mlange, ) have brought attention to energy and cost efficiency of LLM inference. However, optimizing for energy consumption or cost does not directly optimize carbon impact. Carbon optimization goes beyond energy efficiency and cost considerations, requiring a holistic understanding of factors such as the carbon intensity of the energy grid, the lifetime utilization of the host (including CPU, memory, and storage) and the accelerator system, resource provisioning decisions, and the long-term environmental impacts of the underlying infrastructure.
3. Understanding AI Systems’ Carbon Footprint
@@
Figure 3. Trends in bit density (left) and embodied carbon footprint (right) across various DRAM memory technologies for 3 different manufacturers.
@@
Based on recent efforts quantifying the carbon footprint of computing systems (ACT, ; SCARIF, ; eeckhout2024focal, ), the aggregate carbon emissions of AI systems is the combination of operational and embodied carbon as shown in the equation below:
$CF_{Task}=CF^{emb}+CF^{op}=(P_{host}+P_{gpu})\times t\times CI+CF^{emb}_{host}\frac{t}{LT}+CF^{emb}_{gpu}\frac{t}{LT}$
@@
where $P$ represents the power consumption of the host and GPU, $t$ represents the workload run-time, $LT$ represents the hardware lifetime, and $CF^{emb}$ represents the host and GPU embodied carbon. In terms of operational emissions, we measure the power consumption of host-processing systems including CPUs, DRAM, SSD, using a combination of Intel RAPL (RAPL, ); the power consumption of GPUs is empirically measured using NVML (nvml, ). The carbon intensity of the different data center power grids is based on geo-temporal variance of power grid carbon intensities (acun2023carbon, ; greensku, ; watttime, ; electricitymaps, ). To estimate embodied carbon, we leverage a combination of product environmental reports and life cycle analyses from industry (dellr740, ) and open-source tools to quantify the emissions from hardware manufacturing including the Architectural Carbon Modeling Tool (ACT), SCARIF, and GreenSKU (ACT, ; SCARIF, ; greensku, ). There is a void in fine-grained modeling of different memory technology nodes for GPUs, which significantly impacts the embodied carbon footprint estimation. Furthermore, the peripheral components for cooling or power delivery are also not considered in the past models, which is essential for higher-end GPUs (ACT, ; SCARIF, ).
In this section we describe how we extend current methods to study the embodied carbon of AI inference platforms including host systems, motherboards, and GPUs.
3.1. Modeling Embodied Carbon of AI Systems
@@
The embodied carbon footprint (CF) owes to a combination of components for the host systems and GPUs. For host systems, we must account for the main CPU die, DRAM memory, SSD and HDD storage, printed wiring board (PWB), power delivery network (PDN), and server chassis. Similarly, for the GPUs we must account for SoC, DRAM memory, PDN, cooling (e.g., heat sink), and PWB. Table 1 describes the salient parameters in the embodied carbon model for AI inference systems. For many components including PDN and cooling (e.g., heat sink) given the lack of specific technology information we rely on publicly available product environmental reports from Dell (dellr740, ). Below we detail the methodology for estimating the embodied carbon of the remaining salient components.
3.1.1. Integrated circuits (Application Processors)
@@
To compute the embodied carbon emissions of application processors (e.g., CPU die, GPU die) we rely on recently published architectural carbon modeling tools (ACT, ; SCARIF, ; imec-dtco20, ; eeckhout2024focal, ). For application processors we input the process technology and die area into ACT (ACT, ; imec-dtco20, ).
3.1.2. Memory and Storage